Borner Communications’ articles of incorporation authorized the issuance of 125 million common shares. The transactions described below effected changes in Borner’s outstanding shares. Prior to the transactions, Borner’s shareholders’ equity included the following:

Shareholders’ Equity ($ in millions)
Common stock, 105 million shares at $1 par $ 105
Paid-in capital – excess of par 315
Retained earnings 295

Required:
Assuming that Borner Communications retires shares it reacquires (restores their status to that of authorized but unissued shares), record the appropriate journal entry for each of the following transactions: (If no entry is required for a transaction/event, select "No journal entry required" in the first account field. Enter your answers in millions (i.e., 10,000,000 should be entered as 10).)

  1. On January 7, 2021, Borner reacquired 4 million shares at $5.00 per share.
  2. On August 23, 2021, Borner reacquired 6 million shares at $3.50 per share.
  3. On July 25, 2022, Borner sold 5 million common shares at $5 per share.
Answer #1
General Journal Debit Credit
January 7, 2021 Common stock 4
Paid-in capital-excess of par 12 =4*3
Retained earnings 4
     Cash 20 =4*5
August 23, 2021 Common stock 6
Paid-in capital-excess of par 18 =6*3
     Paid in Capital-Share repurchase 3
     Cash 21 =6*3.50
July 25, 2022 Cash 25 =5*5
     Common Stock 5
     Paid-in capital-excess of par 20